# Better Language Models Exhibit Higher Visual Alignment

### v2

@@ Abstract
How well do text-only Large Language Models (LLMs) naturally align with the visual world? We provide the first direct analysis by utilizing frozen text representations in a discriminative vision-language model framework and measuring zero-shot generalization on unseen classes. We find decoder-based LLMs exhibit high intrinsic visual alignment. In particular, more capable LLMs reliably demonstrate stronger generalization. Moreover, utilizing frozen LLMs leads to strong gains in cross-lingual settings, where our approach surpasses CLIP’s accuracy of with for Chinese. Our proposed method improves both robustness and generalization and also significantly reduces the need for paired data and compute, making vision-language models more accessible and adaptable. Code and models are published on our Project Page
Better Language Models Exhibit Higher Visual Alignment
Jona Ruthardt FunAI Lab University of Technology Nuremberg Gertjan J. Burghouts Intelligent Imaging TNO
Serge Belongie Department of Computer Science University of Copenhagen Yuki M. Asano FunAI Lab University of Technology Nuremberg
@@ 1 Introduction
Large Language Models (LLMs) are solely pretrained on unimodal textual data, yet are increasingly incorporated into systems that perceive and interact with the natural world (Ahn et al., 2022; Driess et al., 2023; Wayve, 2023). The lack of direct sensory experience raises fundamental questions as to what extent such models can generalize across modalities and develop a meaningful and accurate understanding of visual reality. Do these models merely regurgitate visually relevant factual knowledge from their training corpus, or do they form internal representations that correspond to real-world phenomena? Despite their successful integration into large-scale Vision-Language Models (VLMs), judging the visual capabilities already inherent to LLMs is difficult. This is not only because of widely varying training recipes and proprietary data sources but particularly due to fine-tuning with paired image-text data, which intertwines with any visual knowledge already contained in text-only models.
@@
In contrast, Sharma et al. (2024) and Huh et al. (2024) more immediately assess the visual nature of LLMs and highlight a non-trivial degree of visual understanding and cross-modal alignment. These works compile proxy tasks or measures such as generating code to represent real-world concepts (Sharma et al., 2024) or correlating visual- with language-based representations (Huh et al., 2024). However, reliance on highly constrained and synthetic tasks with limited practical significance fails to gauge the aptitude of LLMs when deployed in more realistic settings.
@@
To this end, we assess visual alignment—the degree to which language model representations structurally and semantically correspond to those of vision models—through the task of zero-shot open-vocabulary image classification, as popularized by CLIP (Radford et al., 2021). This involves learning a projection of language embeddings into the latent space of a vision model and selecting the label whose representation is most similar to a given image. To ensure a rigorous evaluation of true zero-shot generalization, we enforce strict disjointness between seen and unseen concepts (Lampert et al., 2009). This mitigates concept leakage, a common issue in VLMs trained on web-scale data. For example, Xu et al. (2024) find significant overlap between CLIP’s training data and common evaluation datasets and other works (Fang et al., 2022; Udandarao et al., 2024; Parashar et al., 2024; Mayilvahanan et al., 2025) demonstrate and sharp drops in recognition performance for less-frequent or truly novel concepts. In our proposed setup, generalization relies solely on the semantic information and visual knowledge encoded in language representations. By probing how well models capture visual semantics, we provide insight into their ability to structure and encode text for vision-language applications.
@@
We test a range of language model types and find that features extracted from large modern decoder-based LLMs are more effective than classic encoder-based embeddings, like BERT. Intriguingly, we find that general LLM capability, as measured by MMLU-Pro (Hendrycks et al., 2021b), correlates positively with the model’s visual performance, as shown in Figure 1. Even off-the-shelf LLMs without embedding-specific fine-tuning exhibit strong visual representation abilities.
@@
Finally, we integrate frozen vision and language representations into a lightweight discriminative VLM, ShareLock, demonstrating strong multimodal capabilities across a range of tasks. By capitalizing on the broad pretraining of modern LLMs, ShareLock achieves remarkable cross-lingual zero-shot generalization to non-English languages, outperforming CLIP’s performance of with for Chinese. Despite using only a fraction of the data and learnable parameters, our method approaches the performance of fully optimized CLIP models trained on orders of magnitude more paired data. The visual and linguistic expressiveness of LLM representations is particularly effective in nuanced and fine-grained tasks, resulting in above-CLIP compositional reasoning performance. Our results highlight the considerable extent to which language representations inherently capture visual structure and semantics, despite their unimodal pretraining.
@@
We systematically assess the visual alignment inherent to language models through the lens of strict zero-shot image classification.
Our analysis highlights modern LLMs as effective sources of visual knowledge, with semantically meaningful representations extractable from their internal states.
With ShareLock, we incorporate frozen LLMs with intrinsic visual alignment into VLMs, resulting in improved robustness and generalization that we demonstrate on various tasks involving classification, retrieval, multi-lingual understanding, and compositional reasoning.
@@ 3.1 Zero-Shot Generalization
In order to quantitatively gauge the visual aptitude of language models, we draw on traditional zero-shot learning methodology (cf. Lampert et al. (2009)) and assess how their representations facilitate generalization to novel concepts. To rigorously assess true generalization performance without concept leakage from supervision with arbitrary web-scraped image-captions pairs, we split image classification datasets into seen classes (training stage) and unseen classes (testing stage) with . In the absence of image-specific captions, text-based class representations are used as supervision signals during training and for zero-shot transfer during inference. Crucially, this implies that the performance to discriminate novel concepts is contingent on the validity and cross-modal continuity of the class representations. Therefore, this setup allows us to assess the degree to which language models encode visual knowledge and semantics.
@@ 3.2 Shared Vision-Language-Locked Tuning
To map textual inputs into visual latent spaces, we draw inspiration from late-fusion architectures in CLIP-like models. Texts are first encoded using a language model and subsequently projected into the -dimensional latent space of the vision encoder via a learnable projection network . The latent representation for a given input image or caption is therefore computed by and , respectively. After normalization, their similarity is computed as the cosine similarity, given by the dot product of the embeddings.
@@ 3.3 Experimental Setup
For a comprehensive evaluation, we select four datasets: AWA2 (Xian et al., 2017), CUB (Wah et al., 2011), FGVCAircraft (Maji et al., 2013), and ImageNet+, spanning natural and human-made artifacts, coarse and fine-grained categories, and varying scales (). ImageNet+ considers ImageNet-1k (Deng et al., 2009) classes as seen concepts and uses the 500 most populated ImageNet-21k classes as unseen ones. For AWA2 and CUB, we use splits by Xian et al. (2017) while randomly dividing aircraft types into 50 seen and 20 unseen classes. We report the average per-class classification accuracy across the datasets as a measure of visual generalization ability facilitated by the language embeddings.
@@ 4 Language-driven Visual Generalization
The class-wise supervision and limited concept diversity of conventional image classification datasets can impede vision-language alignment but permit more sophisticated semantic class representations beyond simple template-based targets as typically used with CLIP-like models (Radford et al., 2021; LAION AI, 2022). We thus examine how the nature and information content of different textual class representations111Details about the characteristics and acquisition of these class representations are elaborated in Section B impact generalization performance. The results are summarized in Table 1.
@@
| Language Model | Names | Wikip. | Wikip. | |
| Enc. | BERT-Large (Devlin et al., 2019) | |||
| ModernBERT (Nussbaum et al., 2024) | ||||
| all-roberta-v1 (Liu et al., 2019) | ||||
| T5-XL (Raffel et al., 2020) | ||||
| SentenceT5-XXL (Ni et al., 2022) | ||||
| Flan-UL2 (10B) (Tay, 2024) | ||||
| Dec. | Gemma-2 9B (Mesnard et al., 2024) | |||
| Llama-3 8B (Dubey et al., 2024) | ||||
| NV-Embed-v2 (Lee et al., 2024) |
In line with previous studies (Pratt et al., 2022; Menon and Vondrick, 2023; Yang et al., 2022; Saha et al., 2024), we find that the addition of auxiliary information, such as Wikipedia articles, results in improved performance for most language models. We also find that LLM-generated articles describing a class in the style of Wikipedia (“LLM Wikip.” in Table 1) can provide strong targets during multi-modal alignment, achieving the best overall performance of . Interestingly, relying on strictly human-curated data in the form of actual Wikipedia articles tends only to provide marginal benefits, for example, from and , for NV-Embed and Llama-3. Thus, LLMs can effectively absorb and interpolate substantial amounts of factual information from their training data, positioning them as valuable sources of visually relevant knowledge.
@@
A new insight resulting from this analysis is the competitiveness of decoder-based language models for representing visual concepts. Compared to encoders, we find that representing inputs with decoders can result in higher performance for visual tasks, mirroring a recently emerging trend in the language domain (Lee et al., 2024; Springer et al., 2024). This is likely because these LLMs are trained with vast (and proprietary) training data and are typically much larger than common encoder LMs. In particular, Gemma-2 9B (Mesnard et al., 2024) emerges as the best performer across various types of input data with a maximum performance of . The best-performing encoder, Flan-UL2 (Tay, 2024), only achieves a score of on the same input data, with T5- and particularly BERT-based models exhibiting significantly lower performances. In contrast, other LLMs like Llama or NV-Embed mirror Gemma-2’s performance level more closely.
@@
In Figure 1, we compare various LLMs by the visual generalization ability they possess, as well as their MMLU-Pro (Wang et al., 2024b) score taken from the Open LLM Leaderboard (Fourrier et al., 2024), a common metric to measure LLM performance. We find that the general capability of language models is strongly correlated with their ability to perform well on the visual tasks (Pearson coefficient : ). Within and across model families, we see improved visual generalization as the capacity and capabilities of models increase. Since models steadily improve in the language domain, our evaluation protocol will be helpful in assessing whether the trend of increasing visual understanding will continue in future LLM models. If this holds, VLMs that incorporate LLMs can piggyback off and benefit from developments in the language modeling domain.

### v3

@@ Abstract
How well do text-only large language models (LLMs) align with the visual world? We present a systematic evaluation of this question by incorporating frozen representations of various language models into a discriminative vision-language framework and measuring zero-shot generalization to novel concepts. We find that decoder-based models exhibit stronger visual alignment than encoders, even when controlling for model and dataset size. Moreover, language modeling performance correlates with visual generalization, suggesting that advances in unimodal LLMs can simultaneously improve vision models. Leveraging these insights, we propose ShareLock, a lightweight method for fusing frozen vision and language backbones. ShareLock achieves robust performance across tasks while drastically reducing the need for paired data and compute. With just 563k image-caption pairs and under one GPU-hour of training, it reaches 51% accuracy on ImageNet. In cross-lingual settings, ShareLock dramatically outperforms CLIP, achieving 38.7% top-1 accuracy on Chinese image classification versus CLIP’s 1.4%. Code is available.
@@ 1 Introduction
Large Language Models (LLMs) are solely pretrained on unimodal textual data, yet are increasingly incorporated into systems that perceive and interact with the natural world (Ahn et al., 2022; Driess et al., 2023; Wayve, 2023). The lack of direct sensory experience raises fundamental questions as to what extent such models generalize across modalities and develop a meaningful understanding of visual reality. Do these models merely regurgitate visually relevant factual knowledge from their training corpus, or do they form internal representations that correspond to real-world phenomena? Despite their successful integration into large-scale Vision-Language Models (VLMs), judging the visual capabilities already inherent to LLMs is difficult. This stems not only from differences in training recipes and proprietary data, but especially from fine-tuning with paired image–text data, which blends with the visual knowledge already embedded in text-only models.
@@
In contrast, Sharma et al. (2024) and Huh et al. (2024) more immediately assess the visual nature of LLMs and highlight a non-trivial degree of visual understanding and cross-modal alignment. These works compile proxy tasks such as generating code to represent real-world concepts (Sharma et al., 2024) or correlating vision and language features (Huh et al., 2024). However, reliance on highly constrained and synthetic tasks with limited practical significance fails to gauge the aptitude of LLMs in more realistic settings.
@@
To this end, we assess visual alignment—the degree to which language model representations structurally and semantically correspond to those of vision models—through the task of zero-shot open-vocabulary image classification, as popularized by CLIP (Radford et al., 2021). This involves learning a projection of language embeddings into the vision manifold and selecting the label whose representation is most similar to a given image. To ensure a rigorous evaluation of true zero-shot generalization, we enforce strict disjointness between concepts encountered during training and testing as illustrated in Figure 2 (center) (Lampert et al., 2009). This mitigates concept leakage, a common issue in VLMs (Fig. 2, left). For example, Xu et al. (2024) find significant conceptual overlap between CLIP’s training and evaluation datasets. Other works demonstrate sharp performance drops for less-frequent or truly novel concepts (Fang et al., 2022; Udandarao et al., 2024; Parashar et al., 2024; Mayilvahanan et al., 2025). In our proposed setup, generalization relies solely on the semantic information and visual knowledge encoded in language representations. By probing how well models capture visual semantics, we provide insight into their ability to encode text for vision-language applications.
@@
We observe non-trivial zero-shot generalization across language model types, indicating latent visual-semantic alignment. In particular, features from modern generative LLMs outperform classic encoder-based embeddings such as BERT (Devlin et al., 2019). This advantage of decoder-based models persists when controlling for pretraining data and parameter count across architectures. Intriguingly, we find that general LLM capability, as measured by MMLU-Pro (Hendrycks et al., 2021a), correlates positively with the model’s visual performance, as shown in Figure 1. Even off-the-shelf LLMs without embedding-specific fine-tuning exhibit strong visual representation abilities.
@@
Finally, we integrate frozen vision and language representations into a lightweight discriminative VLM, ShareLock, demonstrating strong multimodal capabilities across a range of tasks. Despite using only a fraction of the data and learnable parameters, our method approaches the performance of fully optimized CLIP models trained on orders of magnitude more paired data. By capitalizing on the broad pretraining of modern LLMs, ShareLock achieves remarkable cross-lingual zero-shot generalization to non-English languages, outperforming CLIP’s performance of with for Chinese. The visual and linguistic expressiveness of LLM representations is particularly effective in nuanced and fine-grained tasks, resulting in above-CLIP compositional reasoning performance. Our results highlight the considerable extent to which language representations capture visual structure and semantics, enabling highly efficient alignment with vision embeddings using limited supervision and parameterization.
@@
We provide a systematic evaluation of the visual alignment inherent to language models, using strict zero-shot image classification as a practically-relevant probing task.
Our analysis highlights modern decoder-based LLMs as effective sources of visual knowledge, with semantically meaningful representations extractable from their internal states.
With ShareLock, we incorporate frozen LLMs with high intrinsic visual alignment into a lightweight VLM, resulting in improved robustness and generalization on various classification, retrieval, multi-lingual understanding, and compositional reasoning tasks.
@@ 3 Probing LLMs for Visual Alignment
This section details our methodology for quantitatively assessing the visual alignment of language models. We introduce the zero-shot evaluation protocol, describe the architecture aligning frozen vision and language backbones, and summarize implementation details to ensure reproducibility.
@@ 3.1 Zero-Shot Generalization
We determine the visual aptitude of language models by drawing on traditional zero-shot learning methodology (cf. Lampert et al. (2009)) and consider how their representations facilitate generalization to novel concepts. To rigorously assess true generalization performance without concept leakage from supervision with arbitrary web-scraped image-captions pairs, we split image classification datasets into aligned classes (training stage) and unaligned classes (testing stage) with (see Figure 2, center). In the absence of image-specific captions, text-based class representations are used as supervision signals during training and for zero-shot transfer during inference. Crucially, this implies that the performance to discriminate novel concepts is contingent on the validity and cross-modal continuity of the class representations. Therefore, this setup allows us to assess the degree to which language models encode visual knowledge and semantics.
@@ 3.2 Shared Vision-Language-Locked Tuning
To map textual inputs into visual latent spaces, we draw inspiration from late-fusion architectures in CLIP-like models. Texts are first encoded using a language model and subsequently projected into the -dimensional latent space of the vision encoder via a learnable projection network . The latent representation for a given input image or caption is therefore computed by and , respectively. Their similarity is computed as the cosine similarity, given by the dot product of the normalized embeddings.
@@ 3.3 Experimental Setup
For a comprehensive evaluation, we select four datasets: AWA2 (Xian et al., 2017), CUB (Wah et al., 2011), FGVCAircraft (Maji et al., 2013), and ImageNet+, spanning natural and human-made artifacts, coarse and fine-grained categories, and varying scales (). ImageNet+ treats ImageNet-1k (Deng et al., 2009) classes as aligned concepts and the 500 most populated ImageNet-21k classes as unaligned ones. For AWA2 and CUB, we use splits by Xian et al. (2017) while randomly dividing aircraft types into 50 aligned and 20 unaligned classes. We report the average per-class classification accuracy on unaligned classes across the datasets as a measure of visual generalization ability facilitated by the language embeddings.
@@ 4 Language-driven Visual Generalization
The class-wise supervision and limited concept diversity of conventional image classification datasets can impede vision-language alignment but permit more sophisticated semantic class representations beyond simple template-based targets as typically used with CLIP-like models (Radford et al., 2021; LAION AI, 2022). We thus examine how the nature and information content of different textual class representations222Details about the characteristics and acquisition of these class representations are elaborated in Appendix C impact generalization performance. The results are summarized in Table 1. As a naive language-free baseline, we also construct one-hot-encoded class representations. Unsurprisingly, the lack of semantic continuity connecting aligned and unaligned concepts results in a near-random generalization score of and demonstrates that generalization in the proposed evaluation protocol is fundamentally dependent on the semantic alignment of vision and language embeddings.
@@
| Language Model [# Parameters] | Names | Wikip. | Wikip. | |
| Enc. | BERT-Large [336M] (Devlin et al., 2019) | |||
| ModernBERT [395M] (Nussbaum et al., 2024) | ||||
| all-roberta-v1 [355M] (Liu et al., 2019) | ||||
| SentenceT5-XL [1B] (Ni et al., 2022) | ||||
| SentenceT5-XXL [5B] (Ni et al., 2022) | ||||
| Flan-UL2 [10B] (Tay, 2024) | ||||
| Dec. | NV-Embed-v2 [8B] (Lee et al., 2024) | |||
| Llama-3 [8B] (Dubey et al., 2024) | ||||
| Gemma-2 [9B] (Mesnard et al., 2024) |
We find that the addition of auxiliary information, such as Wikipedia articles, results in improved performance for most language models, reflecting insights from previous studies (Pratt et al., 2022; Menon and Vondrick, 2023; Yang et al., 2022; Saha et al., 2024). We also find that LLM-generated articles describing a class in the style of Wikipedia (“LLM Wikip.” in Table 1) can provide strong targets during multi-modal alignment, achieving the best overall performance of . Interestingly, relying on strictly human-curated data in the form of actual Wikipedia articles tends only to provide marginal benefits, for example, from and for Llama-3 and NV-Embed. Thus, LLMs can effectively absorb and interpolate substantial amounts of factual information from their training data, positioning them as valuable sources of visually relevant knowledge.
@@
A new insight resulting from our analysis is the competitiveness of decoder-based language models in representing visual concepts. Compared to language encoders commonly used in vision tasks, we find that representing inputs with decoders can result in higher performance, mirroring a recently emerging trend in the language domain (Lee et al., 2024; Springer et al., 2024). However, results in Table 1 confound factors such as model and dataset size with the different pretraining tasks. To isolate the impact of architecture, we conducted controlled comparisons with the Ettin model family (Weller et al., 2025) that only varies attention patterns and training objectives while unifying data and parameter count for encoder- and decoder-based models. Figure 3 reveals that decoders outperform encoders by an average of percentage points across model sizes, suggesting that autoregressive next-token prediction gives rise to more vision-aligned representations compared to masked language modeling in encoder models. Beyond this intrinsic advantage, decoder models are often pretrained on larger-scale data and comprise more parameters, leading to more capable models in addition to architectural benefits. For instance, Gemma-2 9B achieves the highest score of across our evaluated input types, outperforming the strongest and similarly sized encoder model, Flan-UL2, which reaches on the same input data. Other decoders like LLaMA and NV-Embed perform on par with Gemma, whereas encoder models such as T5- and BERT-based variants lag considerably behind. These findings suggest that decoder-based LLMs can offer both architectural and practical advantages for visual representation and alignment tasks.
@@
In Figure 1, we compare various LLMs by the visual generalization ability they possess, as well as their MMLU-Pro (Wang et al., 2024a) score taken from the Open LLM Leaderboard (Fourrier et al., 2024), a common metric to measure LLM performance. We find that the general capability of language models is strongly correlated with their ability to perform well on the visual tasks (Pearson coefficient : ). Within and across model families, we see improved visual generalization as the capacity and capabilities of models increase. Since models steadily improve in the language domain, our evaluation protocol will be helpful in assessing whether the trend of increasing visual understanding will continue in future LLM models. If this holds, VLMs that incorporate LLMs can piggyback off developments in the language modeling domain.
5 LLMs for General-Purpose VLMs
The previous section demonstrated that LLMs trained for textual next-token prediction encode visually-aligned semantics in their representations. Next, we investigate whether this inherent visual alignment can be leveraged to obtain effective and efficient multimodal models by integrating off-the-shelf LLMs directly into CLIP-like VLMs. For this, we relax the strict zero-shot setup and utilize larger-scale image-caption datasets (see Figure 2, left) to explore the benefits and limitations of general-purpose VLMs that fuse existing frozen foundation models with minimal data and compute.
5.1 Experimental Setup
Methodology.
We propose integrating LLMs into a CLIP-like framework through “Shared Vision-Language-Locked Tuning“ (ShareLock), extending the architecture and training setup introduced in Section 3. As illustrated in Figure 4, this design places a LLM at the core of the model and maps its expressive representations into the feature space of a frozen vision model. If not noted otherwise, Llama-3 8B (Dubey et al., 2024) and DINOv2-L (Oquab et al., 2023) are used as frozen backbones. Given the greater availability of training data compared to the strict zero-shot setting, we expand the projection network applied on top of the frozen language features to four layers with a hidden dimension of . The additional parameters, combined with ReLU activations between layers, enhance ShareLock’s ability to capture complex, non-linear cross-modal correspondences. Additionally, the maximum number of optimization steps is increased to 5000.
Datasets.
Our investigation focuses on leveraging LLMs with minimal additional paired data and explores how unimodal embeddings can drive robust multimodal performance with minimal supervision and alignment. As a result, our evaluation is limited to comparably small paired datasets. COCO Captions. COCO Captions (Chen et al., 2015) contains 83k images with multiple human-written captions per image, from which we randomly sample during training. CC3M. Conceptual Captions (Sharma et al., 2018) comprises 2.8M filtered web-scraped image-alt-text pairs. We also utilize a smaller subset with more balanced concept coverage designed for LLaVA (Liu et al., 2023). CC12M. CC12M (Changpinyo et al., 2021) expands CC3M by 12M image-text pairs, enabling evaluation at larger scales. Due to expired links, our version contains about 8.5M samples. If not noted otherwise, CC3M is used to compare design choices.
Computational efficiency and storage.
Precomputing features for the CC3M-Llava subset (563k pairs) using Llama-3 8B requires around 8 hours on a single A100 40GB GPU. Extracting DINOv2 features and optimizing the MLP-based projection network take approximately 1 GPU hour each, totaling roughly 10 GPU hours. This is paired with a significant storage reduction from over 80GB for the raw data to just 12GB for precomputed features. For classification at inference, language targets are computed once per label set and inference cost is dominated by the vision encoder. Consequently, CLIP, LiT, and ShareLock have comparable FLOPs and latency when using the same backbone architecture (e.g., ViT-L/14@224x224: GFLOPs), despite ShareLock’s larger LLM.
Evaluation.
We employ a comprehensive suite of evaluations to assess ShareLock’s capabilities across a wide range of tasks. Based on the publicly available CLIP Benchmark (LAION AI, 2022), we gauge the models’ zero-shot classification abilities across diverse datasets: ImageNet-1k (Deng et al., 2009), ImageNet-R(endition) (Hendrycks et al., 2021b), ImageNet-A(dversarial) (Hendrycks et al., 2021c), ImageNet-S(ketch) (Wang et al., 2019), Oxford-IIIT Pets (Parkhi et al., 2012), Oxford Flowers (Nilsback and Zisserman, 2008), Stanford Cars (Krause et al., 2013), and FGVCAircraft (Maji et al., 2013). We also provide qualitative text-to-image retrieval results on ImageNet for CC3M-trained models. Moreover, the challenging compositionality Winoground task (Thrush et al., 2022) is explored.
Baselines and comparisons.
We compare our straightforward and economical way of incorporating LLMs into VLMs to more conventional CLIP-like methods, particularly emphasizing data-efficient alignment approaches. Alongside the original ViT-B/16 variant of CLIP (Radford et al., 2021), we test against several CLIP-like models trained on public datasets of different scales and with modified learning objectives that all share the standard dual-encoder architecture where both image and text encoders are trained from scratch with a contrastive loss (Fig. 4, left) (Fan et al., 2023; Gadre et al., 2023; Mu et al., 2022). Leveraging pretrained models, we evaluate how ShareLock compares to LiT (Zhai et al., 2022) and ASIF (Norelli et al., 2023) by reproducing these methods on smaller-scale datasets. LiT corresponds to the partially frozen setup in Figure 4 (center), where the vision encoder is locked but the text tower and projection are fine-tuned. ShareLock follows the fully frozen setup in Figure 4 (right), freezing both vision and language encoders and learning only lightweight projection networks. For LiT baselines, we initialize the language encoder with pretrained BERT-Base weights (Devlin et al., 2019), following Zhai et al. (2022). When comparing ShareLock with LiT and ASIF, the same precomputed features (except for LiT’s language inputs) are used.
5.2 Comparison to Conventional VLMs